OpenStudy (ineedhelpnow):

if 10 pens and 15 pads of paper cost $26, and 5 of the same pens and 10 of the same pads cost $16, find the cost of a pen and a pad.

OpenStudy (anonymous):

\[10x+15y=26\] \[5x+10y=16\] is what you need to solve

OpenStudy (anonymous):

multiply second equation by -2 to get \[10x+15y=26\] \[-10x-20y=-32\] add to get \[-5y=-6\] \[y=\frac{6}{5}=1.2\]

OpenStudy (anonymous):

so pads cost $1.20

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

What about pen?

OpenStudy (akshay_budhkar):

find x..

OpenStudy (anonymous):

we substitute y = 1.2 into equation to solve for x

OpenStudy (akshay_budhkar):

right

OpenStudy (anonymous):

well if you know the price of the pad, you certainly know the price of the pen. ten pads cost $12 so 5 pens cost $4 so each pen is $.80

OpenStudy (ineedhelpnow):

okay. just wanted to check my answer. i made the mistake of dividing -6 by something to get .83, but i get it now.

OpenStudy (anonymous):

\[\checkmark\]

OpenStudy (akshay_budhkar):

how to do that satellite?